OpenStudy (anonymous):

Consider the line that passes through the point (-1, 6) and has a slope of -4. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (anonymous):

But no dont show ALL work. Just help me understand it :) thanks

OpenStudy (apoorvk):

The equation of a line passing through \((x_1, y_1)\) and having a slope \(m\) is: \(\large \frac{y - y_1}{x - x_1} = m\) can you use this?

OpenStudy (anonymous):

Part 1 1) use point -slope form y-y1=m(x-x1) so the equation would look like y-6=-4(x+1) Part Two 2) then the equation becomes y-6=-4x-4 3) add six to both sides to get y=-4x+2 Part 3 4) use the slope and y-intercept of the above equation 5) the standard from would look like y-2=x+0